# MIFNet: Learning Modality-Invariant Features for Generalizable Multimodal Image Matching

### v1

@@ Abstract
Many keypoint detection and description methods have been proposed for image matching or registration. While these methods demonstrate promising performance for single-modality image matching, they often struggle with multimodal data because the descriptors trained on single-modality data tend to lack robustness against the non-linear variations present in multimodal data. Extending such methods to multimodal image matching often requires well-aligned multimodal data to learn modality-invariant descriptors. However, acquiring such data is often costly and impractical in many real-world scenarios. To address this challenge, we propose a modality-invariant feature learning network (MIFNet) to compute modality-invariant features for keypoint descriptions in multimodal image matching using only single-modality training data. Specifically, we propose a novel latent feature aggregation module and a cumulative hybrid aggregation module to enhance the base keypoint descriptors trained on single-modality data by leveraging pre-trained features from Stable Diffusion models. We validate our method with recent keypoint detection and description methods in three multimodal retinal image datasets (CF-FA, CF-OCT, EMA-OCTA) and two remote sensing datasets (Optical-SAR and Optical-NIR). Extensive experiments demonstrate that the proposed MIFNet is able to learn modality-invariant feature for multimodal image matching without accessing the targeted modality and has good zero-shot generalization ability. The source code will be made publicly available.
@@ I Introduction
Multimodal image matching is essential in various domains, including medical imaging, remote sensing, and autonomous driving. Its primary goal is to align images captured by different imaging sensors, thereby providing complementary information for downstream tasks such as retinal disease monitoring [1], laser eye surgeries [2], image fusion [3], change detection [4, 5], and image mosaicking [6]. However, multimodal image matching faces significant challenges due to the inherent differences between modalities. A major challenge lies in the nonlinear intensity variations inherent between multimodal images. Additionally, geometric distortions and viewpoint discrepancies caused by different imaging conditions further complicate the matching process.
@@
To effectively utilize latent diffusion features, we first refine the coarse latent diffusion features before integrating them with the base features. Specifically, we introduce a novel Latent Feature Aggregation (LFA) module, which employs a Gaussian mixture model to refine coarse latent features, enhancing both semantic and modality invariance. Subsequently, we design a Cumulative Hybrid Aggregation (CHA) module to fuse the refined features with base features using multiple self-attention and cross-attention layers, generating the final modality-invariant features. Both modules are trained in a self-supervised manner by comparing training samples with their randomly augmented counterparts, eliminating the need for additional annotations. Remarkably, this process enables the learning of modality-invariant features using only single-modality training images. As shown in Fig. 1, MIFNet demonstrates strong modality invariance, significantly improving the cross-modality matching performance of state-of-the-art methods (e.g., XFeat [20]) across diverse datasets.
@@
Figure 2: Overview of our MIFNet framework. (1) Given an image $I_{A}$, we first apply a random transformation to obtain image $I_{B}$. Then, we extract base features ($F_{A}^{b},F_{B}^{b}$) and latent diffusion features ($F_{A}^{l},F_{B}^{l}$) using a frozen keypoint detector and the Stable Diffusion v2 model, respectively. (2) The proposed latent feature aggregation which includes self-attention, cross-attention and Gaussian Mixture Model (GMM) to refine the aggregated latent features. (3) The proposed cumulative hybrid aggregation module integrate refined latent features and base features through multi-layer attention mechanisms to generate modality-invariant features (${F}_{A}^{m},{F}_{B}^{m}$). (4) Finally, the matching module predicts the matching score matrix between features and is supervised by the correspondence between images.
@@ II-B Multimodal Image Matching
Different from these methods, MIFNet requires only single-modality images as the training data to train the models for modality-invariant features, significantly reducing the cost to collect well aligned multimodal data. Moreover, MIFNet demonstrates excellent generalization capabilities, achieving strong multimodal image matching performances across various unseen modality datasets.
@@ III-A Original Feature Extraction
Latent Diffusion Feature Extraction. With advancements in diffusion and foundational models [51, 29, 54], various methods have been developed to extract features with strong cross-domain generalization capabilities. In this paper, we select Stable Diffusion v2 [29] as our feature extractor thanks to its training on billions of image-text pairs, which grants it powerful semantic representation abilities and cross-modal semantic invariance. The diffusion process in this model comprises forward and reverse stages. In the forward diffusion stage, noise is progressively added to clean images, transforming them from a coherent state to a noisy one. In contrast, the reverse diffusion process reconstructs the original image from the noisy state. To retain the model’s intrinsic capabilities, we extract features directly from the frozen decoder layer of the Stable Diffusion v2 during the forward diffusion process. Specifically, we use intermediate features from the second layer of the decoder as the latent feature map. Finally, the latent diffusion features $F_{A}^{l}$ and $F_{B}^{l}$ are sampled from this latent feature map using the normalized coordinates of keypoints extracted by the base feature detector.
@@ III-B Latent Feature Aggregation
The GMM model parameters are optimized by maximizing the log-likelihood function, which is given by:
$\mathcal{L}=\sum_{n=1}^{N}\log\left(\sum_{k=1}^{K}\pi_{k}\mathcal{N}(\mathbf{x}_{n}\mid\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k})\right),$ (4)
where $\mathbf{x}_{n}$ denotes the $n$-th feature point, and $N$ represents the total number of feature points extracted from a single image.
@@
The optimization objective for our proposed LFA is defined:
$\mathcal{L}_{\text{LFA}}=\mathcal{L}_{\text{intra}}-\mathcal{L}_{\text{inter}}.$ (7)
@@ III-D Matching Optimization
The final overall loss $\mathcal{L}$ takes into account both the latent feature aggregation loss and the matching loss:
$\mathcal{L}=\mathcal{L}_{match}+\lambda*\mathcal{L}_{LFA},$ (13)
where $\lambda$ is a hyper-parameter, and is set to 2.0 in all experiments if not otherwise stated.
@@ IV Experiments
Surf-PIFD-RPM [18] T 24.5 43.9 91.2 34.9 56.2 115.6 6.7 48.4 113.2
UR-SIFT-PIFD [42] 27.4 40.1 87.6 29.9 63.4 140.4 6.7 56.9 102.7
KPVSA-NET [25] M 63.9 14.3 37.5 50.4 34.6 81.9 40.0 22.4 47.1
Content-Adaptive [45] 75.7 8.9 24.7 46.9 40.3 85.3 46.7 17.9 38.6
SuperPoint [21] S 73.1 9.7 29.2 17.4 62.0 145.1 6.7 41.3 85.1
@@ IV-B Implementation Details
During MIFNet training, we apply various random data augmentations, including homography transformations and brightness adjustments, to generate synthetic single-modality color image pairs. These pairs are separately processed through a pre-trained frozen keypoint detector and Stable Diffusion v2 [29] to extract keypoints, base features, and diffusion features. Using pixel correspondences between the image pairs, we supervise the training of MIFNet. In the LFA module, the number of clusters $K$ is set to 5. We use the Adam optimizer with a learning rate of $1\times 10^{-4}$, a batch size of 2, and 15 epochs for training. The model is trained on an NVIDIA 3090 GPU. For multimodal image matching evaluation, we use K-nearest neighbors for feature matching and apply cv2.findHomography to compute the homography matrix.
@@ IV-C Zero-shot Evaluation on Multimodal Retinal Image Matching
Experimental Setting. To verify the effectiveness of our method, we train the MIFNet on single-modality color fundus images and test on three challenging datasets under zero-shot setting without any fine-tuning. We use keypoint detection and description methods SuperPoint [21], SuperRetina [36], ALIKED [41], and XFeat [20], as base feature extractors combined with MIFNet to validate its effectiveness. The base detectors are fine-tuned on color fundus training set before being frozen in MIFNet training. We compare our methods ($X$+MIFNet, where $X$ stands for the different base detector) with four different multimodal retinal image matching methods, including traditional methods SURF-PIIFD-RPM [18] and UR-SIFT-PIIFD [42], which do not require training as well as KPVSA-NET [25] and Content-Adaptive [45] which are optimized with domain specific data. We make our best efforts to collect open-source datasets for retraining these network. The datasets include HRF [62], RECOVERY-FA [63], ROSE [64], and MEMO [59].
@@
Figure 4: The visualization results on three multimodal retinal test sets CF-OCT, CF-FA, and EMA-OCTA. The green lines represent correctly matched pairs, while the red lines represent incorrectly matched pairs.
@@
OS-SIFT [17] T 18.6 4.0 3.1 77.7 1.7 34.1
RIFT2 [65] 27.5 4.0 4.9 79.8 1.6 39.2
CMMNet [66] M 22.8 4.3 2.6 80.6 1.5 39.4
ReDFeat [23] 41.4 3.3 9.1 85.4 1.5 46.8
SuperPoint [21] S 15.1 4.4 2.8 74.9 1.9 35.6
@@
On the CF-OCT test set, learning-based single-modality matching algorithms outperform all handcrafted methods [18, 42], with SuperPoint [21] achieving the highest registration success rate of 73.1%. This is due to the primary challenge in the CF-OCT dataset being the scale difference between image pairs, which learning-based methods handle more effectively. For multimodal-specific methods, Content-Adaptive [45] achieves an accuracy of 75.7%, outperforming SuperPoint by converting images into binary vessel masks using a vessel segmentation network. In contrast, KPVSA-NET [25] performs poorly, as its accuracy is constrained by the effectiveness of the CycleGAN network. However, after incorporating our proposed MIFNet, the SuperPoint, SuperRetina, and XFeat methods surpass existing multimodal-specific methods. Notably, MIFNet leads to a 43.6% improvement in the success rate when applied to XFeat [20], achieving the highest success rate of 90.4%.
@@
Figure 5: The visualization results of feature matching in aerial scenes: Optical-SAR and Optical-NIR. The green lines represent correctly matched pairs according to the homography ground truth, while the red lines represent incorrectly matched pairs.
@@ IV-D Zero-shot Evaluation on Multimodal Remote Sensing Image Matching
Experimental Setting. Our method is trained on a single-modality optical dataset and directly tested on two unseen datasets without any fine-tuning. We apply MIFNet to three representative single-modality methods: SuperPoint [22], ALIKED [41], and XFeat [20]. We then compare our approach with two traditional methods, RIFT2 [65] and OS-SIFT [17], which are widely used in multimodal remote sensing image matching, as well as two multimodal methods, CMMNet [66] and ReDFeat [23]. Notably, these methods require paired data specific to each modality pair for training. To facilitate this, we select 2,000 image pairs from the GF3 [60] dataset and 1,000 image pairs from the VEDAI [61] dataset for their training.
@@
We observe that MIFNet enhances cross-modality matching performance for three widely used single-modality image matchers. In the Optical-SAR scenario, which exhibits significant nonlinear intensity differences, MIFNet improves the SRR metric by 24.6%, 21.9%, and 17.7% for SuperPoint, ALIKED, and XFeat, respectively. Notably, SuperPoint combined with MIFNet achieves a registration success rate of 39.7%, only 1.7% behind the best-performing method, ReDFeat [23]. However, ReDFeat is trained on paired multimodal data and lacks generalization to unseen modalities. In contrast, our proposed MIFNet is trained without any exposure to SAR or NIR data.

### v3

@@ Abstract
Many keypoint detection and description methods have been proposed for image matching or registration. While these methods demonstrate promising performance for single-modality image matching, they often struggle with multimodal data because the descriptors trained on single-modality data tend to lack robustness against the non-linear variations present in multimodal data. Extending such methods to multimodal image matching often requires well-aligned multimodal data to learn modality-invariant descriptors. However, acquiring such data is often costly and impractical in many real-world scenarios. To address this challenge, we propose a modality-invariant feature learning network (MIFNet) to compute modality-invariant features for keypoint descriptions in multimodal image matching using only single-modality training data. Specifically, we propose a novel latent feature aggregation module and a cumulative hybrid aggregation module to enhance the base keypoint descriptors trained on single-modality data by leveraging pre-trained features from Stable Diffusion models. We validate our method with recent keypoint detection and description methods in three multimodal retinal image datasets (CF-FA, CF-OCT, EMA-OCTA) and two remote sensing datasets (Optical-SAR and Optical-NIR). Extensive experiments demonstrate that the proposed MIFNet is able to learn modality-invariant feature for multimodal image matching without accessing the targeted modality and has good zero-shot generalization ability. The code will be released at https://github.com/lyp-deeplearning/MIFNet.
@@ I Introduction
Multimodal image matching is essential in various domains, including medical imaging, remote sensing, and autonomous driving. Its primary goal is to register images captured by different imaging sensors, thereby providing complementary information for downstream tasks such as retinal disease monitoring [1], laser eye surgeries [2], image fusion [3], change detection [4, 5], and image mosaicking [6]. However, multimodal image matching faces significant challenges due to the inherent differences between modalities. A major challenge lies in the nonlinear intensity variations inherent between multimodal images. Additionally, geometric distortions and viewpoint discrepancies caused by different imaging conditions further complicate the matching process.
@@
To effectively utilize latent diffusion features, we first refine the coarse latent diffusion features before integrating them with the base features. Specifically, we introduce a novel Latent Feature Aggregation (LFA) module, which employs a Gaussian mixture model to refine coarse latent features, enhancing both semantic and modality invariance. The choice of GMM is motivated by its ability to handle complex, overlapping feature distributions through soft clustering. This allows for better semantic separation and intra-cluster compactness compared to traditional regression models. Subsequently, we design a Cumulative Hybrid Aggregation (CHA) module to fuse the refined features with base features using multiple self-attention and cross-attention layers. The CHA module progressively aggregates features across multiple layers, facilitating both intra-image and cross-image feature interactions, thereby resulting in more precise feature representations. Both modules are trained in a self-supervised manner by comparing training samples with their randomly augmented counterparts, eliminating the need for additional annotations. Remarkably, this process enables the learning of modality-invariant features using only single-modality training images. As shown in Fig. 1, MIFNet demonstrates strong modality invariance, significantly improving the cross-modality matching performance of state-of-the-art methods (e.g., XFeat [20]) across diverse datasets.
@@
Figure 2: Overview of our MIFNet framework. (1) Given an image $I_{A}$, we first apply a random transformation to obtain image $I_{B}$. Then, we extract base features ($F_{A}^{b},F_{B}^{b}$) and latent diffusion features ($F_{A}^{l},F_{B}^{l}$) using a frozen keypoint detector and the Stable Diffusion v2 model, respectively. (2) The proposed latent feature aggregation which includes self-attention, cross-attention and Gaussian Mixture Model (GMM) to refine the aggregated latent features. (3) The proposed cumulative hybrid aggregation module integrates refined latent features and base features through multi-layer attention mechanisms to generate modality-invariant features (${F}_{A}^{m},{F}_{B}^{m}$). (4) Finally, the matching module predicts the matching score matrix between features and is supervised by the correspondence between images.
@@ II-B Multimodal Image Matching
Different from these methods, MIFNet requires only single-modality images as the training data to train the models for modality-invariant features, significantly reducing the cost to collect well registered multimodal data. Moreover, MIFNet demonstrates excellent generalization capabilities, achieving strong multimodal image matching performances across various unseen modality datasets.
@@ III-A Original Feature Extraction
Latent Diffusion Feature Extraction. With advancements in diffusion and foundational models [51, 29, 54], various methods have been developed to extract features with strong cross-domain generalization capabilities. In this paper, we select Stable Diffusion v2 [29] (SDv2) as our feature extractor thanks to its ability to encode rich and transferable semantic information, learned from billions of image-text pairs. Such pretraining enables the model to produce modality-invariant features even without task-specific fine-tuning. The diffusion process in this model comprises forward and reverse stages. In the forward diffusion stage, noise is progressively added to clean images, transforming them from a coherent state to a noisy one. In contrast, the reverse diffusion process reconstructs the original image from the noisy state. To retain the model’s intrinsic capabilities, we extract features directly from the frozen decoder layer of the SDv2 during the forward diffusion process. This design allows us to utilize the pretrained latent space as a plug-and-play semantic prior, preserving both high-level abstraction and spatial structure. Specifically, we use intermediate features from the second layer of the decoder as the latent feature map. Finally, the latent diffusion features $F_{A}^{l}$ and $F_{B}^{l}$ are sampled from this latent feature map using the normalized coordinates of keypoints extracted by the base feature detector. These latent features serve as high-level semantic cues that are less sensitive to low-level appearance variations, thereby enhancing the robustness and discriminability of our framework in challenging cross-modal matching scenarios.
@@ III-B Latent Feature Aggregation
The GMM model parameters are optimized using the Expectation-Maximization (EM) algorithm, which iteratively refines cluster assignments and updates Gaussian parameters to maximize the likelihood of the observed feature distribution.
@@
We employ the combination of the intra-class compactness and inter-class separation objectives to refine the semantic distribution of the latent features.
@@ III-D Matching Optimization
In addition to the matching supervision, we introduce a latent feature aggregation loss to enhance the semantic consistency of the extracted features. The LFA loss is defined as:
$\mathcal{L}_{\text{LFA}}=\mathcal{L}_{\text{intra}}-\mathcal{L}_{\text{inter}},$ (11)
where $\mathcal{L}_{\text{intra}}$ encourages features within the same semantic cluster to be compact, and $\mathcal{L}_{\text{inter}}$ enforces separation between different cluster centers. The final training objective combines the matching loss and the LFA loss as:
$\mathcal{L}=\mathcal{L}_{\text{match}}+\lambda\mathcal{L}_{\text{LFA}},$ (12)
where $\lambda$ is a weighting factor set to 2.0 in all experiments unless otherwise specified.
@@ IV Experiments
Surf-PIFD-RPM (BSPC’15) T 24.5 43.9 91.2 34.9 56.2 115.6 6.7 48.4 113.2
UR-SIFT-PIFD (ICPR’18) 27.4 40.1 87.6 29.9 63.4 140.4 6.7 56.9 102.7
KPVSA-NET (MICCAI’22) M 63.9 14.3 37.5 50.4 34.6 81.9 40.0 22.4 47.1
Content-Adaptive (TIP’21) 75.7 8.9 24.7 46.9 40.3 85.3 46.7 17.9 38.6
Xoftr (CVPR’24) 59.6 17.8 38.5 52.0 29.2 64.3 33.3 20.7 42.3
Superglue (CVPR’20) S 76.9 7.9 16.1 20.8 58.3 127.6 6.7 37.5 64.3
lightglue (ICCV’23) 78.8 7.3 14.6 24.3 52.7 101.1 6.7 33.1 56.5
GIM (DKM) (ICLR’24) 80.8 5.3 12.7 41.8 39.4 90.3 6.7 32.1 59.8
SuperPoint (CVPR’18) 73.1 9.7 29.2 17.4 62.0 145.1 6.7 41.3 85.1
@@ IV-B Implementation Details
During MIFNet training, we randomly select a single image from the training set and apply various random augmentations, such as homography transformations and brightness adjustments, to generate a synthetic training pair consisting of the original and the augmented image. These pairs are separately processed through a pre-trained frozen keypoint detector and Stable Diffusion v2 [29] to extract keypoints, base features, and diffusion features. Using pixel correspondences between the image pairs, we supervise the training of MIFNet. In the LFA module, the number of clusters $K$ for GMM is set to 5. We adopt the standard GaussianMixture implementation from scikit-learn, using K-means for initialization and the Expectation-Maximization (EM) algorithm for parameter optimization. We use the Adam optimizer with a learning rate of $1\times 10^{-4}$, a batch size of 2, and 15 epochs for training. The model is trained on an NVIDIA 3090 GPU. For multimodal image matching evaluation, we use K-nearest neighbors for feature matching and apply cv2.findHomography to compute the homography matrix.
@@ IV-C Zero-shot Evaluation on Multimodal Retinal Image Matching
Experimental Setting. To verify the effectiveness of our method, we train the MIFNet on single-modality color fundus images and test on three challenging datasets under a zero-shot setting without any fine-tuning. We use keypoint detection and description methods including SuperPoint [21], SuperRetina [36], ALIKED [41], and XFeat [20] as base feature extractors combined with MIFNet to validate its effectiveness. The base detectors are fine-tuned on the color fundus training set before being frozen during MIFNet training. We compare our methods ($X$+MIFNet, where $X$ denotes different base detectors) with five multimodal retinal image matching approaches, including traditional handcrafted methods SURF-PIIFD-RPM [18] and UR-SIFT-PIIFD [42], which do not require training, as well as KPVSA-NET [25], Content-Adaptive [45], and Xoftr [62], which are optimized using domain-specific data. We make our best efforts to collect publicly available datasets for retraining. These include HRF [63], RECOVERY-FA [64], ROSE [65], and MEMO [59]. In addition, we also compare against several representative state-of-the-art single-modality matching methods, including SuperGlue [22], LightGlue [33], and GIM [66], to further validate the performance gap under cross-modal settings.
@@
Figure 4: The visualization results on three multimodal retinal test sets CF-OCT, CF-FA, and EMA-OCTA. The green lines represent correctly matched pairs, while the red lines represent incorrectly matched pairs. Each scene includes the results of tests under geometric transformations.
@@
OS-SIFT (TGRS’18) T 18.6 4.0 3.1 77.7 1.7 34.1
RIFT2 (Arxiv’23) 27.5 4.0 4.9 79.8 1.6 39.2
CMMNet (TGRS’21) M 22.8 4.3 2.6 80.6 1.5 39.4
ReDFeat (TIP’23) 41.4 3.3 9.1 85.4 1.5 46.8
Xoftr (CVPR’24) 32.8 3.8 7.7 86.0 1.4 44.9
Superglue (CVPR’20) S 18.0 4.3 3.0 79.0 1.8 38.9
Lightglue (ICCV’23) 19.3 4.3 3.2 81.5 1.7 39.4
GIM (DKM) (ICLR’24) 20.5 4.2 3.8 82.5 1.6 42.0
SuperPoint (CVPR’18) 15.1 4.4 2.8 74.9 1.9 35.6
@@
On the CF-OCT test set, learning-based single-modality matching algorithms outperform all handcrafted methods [18, 42], with GIM [66] achieving the highest registration success rate of 80.8%. This is due to the primary challenge in the CF-OCT dataset being the scale difference between image pairs, which learning-based methods handle more effectively. For multimodal-specific methods, Content-Adaptive [45] achieves an accuracy of 75.7%, outperforming SuperPoint by converting images into binary vessel masks using a vessel segmentation network. In contrast, KPVSA-NET [25] performs poorly, as its accuracy is constrained by the effectiveness of the CycleGAN network. However, after incorporating our proposed MIFNet, the SuperPoint, SuperRetina, and XFeat methods surpass existing multimodal-specific methods. Notably, MIFNet leads to a 43.6% improvement in the success rate when applied to XFeat [20], achieving the highest success rate of 90.4%.
@@
Figure 5: The visualization results of feature matching in aerial scenes: Optical-SAR and Optical-NIR. The green lines represent correctly matched pairs according to the homography ground truth, while the red lines represent incorrectly matched pairs. Each scene includes the results of tests under geometric transformations.
@@ IV-D Zero-shot Evaluation on Multimodal Remote Sensing Image Matching
Experimental Setting. Our method is trained on a single-modality optical dataset and directly tested on two unseen datasets without any fine-tuning. We apply MIFNet to three representative single-modality methods: SuperPoint [22], ALIKED [41], and XFeat [20]. In addition, we also compare our method with several state-of-the-art single-modality matchers, including SuperGlue [22], LightGlue [33], and GIM [66]. We then compare our approach with two traditional methods, RIFT2 [67] and OS-SIFT [17], which are widely used in multimodal remote sensing image matching, as well as two multimodal methods, CMMNet [68], ReDFeat [23] and Xoftr [62]. Notably, these methods require paired data specific to each modality pair for training. To facilitate this, we select 2,000 image pairs from the GF3 [60] dataset and 1,000 image pairs from the VEDAI [61] dataset for their training.
@@
We observe that MIFNet consistently enhances cross-modality matching performance across three widely used single-modality image matchers. In particular, for the Optical-SAR scenario, the matching performance tends to be relatively low for all methods due to substantial nonlinear differences in pixel intensities, severe speckle noise inherent to SAR imaging, and the frequent absence of critical structural details such as road features. Nevertheless, MIFNet significantly improves the Success Rate (SRR) by 24.6%, 21.9%, and 17.7% for SuperPoint, ALIKED, and XFeat, respectively. Notably, the combination of SuperPoint with MIFNet achieves a registration success rate of 39.7%, merely 1.7% lower than the top-performing method, ReDFeat [23]. However, ReDFeat is trained on paired multimodal data and lacks generalization to unseen modalities. In contrast, our proposed MIFNet is trained without any exposure to SAR or NIR data.